OpenStudy (anonymous):

3/7 * 1/3 + 4/7 * 3/8 =Simplify your answer use order of operations

jimthompson5910 (jim_thompson5910):

\[\large \frac{3}{7}*\frac{1}{3}+\frac{4}{7}*\frac{3}{8}\] \[\large \frac{3*1}{7*3}+\frac{4*3}{7*8}\] \[\large \frac{3}{21}+\frac{12}{56}\] \[\large \frac{1}{7}+\frac{3}{14}\] \[\large \frac{1*2}{7*2}+\frac{3}{14}\] \[\large \frac{2}{14}+\frac{3}{14}\] \[\large \frac{2+3}{14}\] \[\large \frac{5}{14}\] ======================================================= So, \[\large \frac{3}{7}*\frac{1}{3}+\frac{4}{7}*\frac{3}{8} = \frac{5}{14}\]

OpenStudy (anonymous):

8/3 divided by (1/4 + 9/16) =

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

i dont know how to divide

jimthompson5910 (jim_thompson5910):

first add 1/4 + 9/16

jimthompson5910 (jim_thompson5910):

what do you get here?

jimthompson5910 (jim_thompson5910):

remember the denominators must be the same

OpenStudy (anonymous):

10/20

jimthompson5910 (jim_thompson5910):

no, the denominators must be the same before you add

OpenStudy (anonymous):

then i dont get it

jimthompson5910 (jim_thompson5910):

1/4 = x/16 what must be x?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so 1/4 = 4/16

jimthompson5910 (jim_thompson5910):

1/4 + 9/16 turns into 4/16 + 9/16

jimthompson5910 (jim_thompson5910):

when you add fractions, you do NOT add the denominators (only the numerators)

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

the denominator will remain 16

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

13/16

jimthompson5910 (jim_thompson5910):

so \[\large \frac{8}{3} \div \left(\frac{1}{4} + \frac{9}{16}\right)\] turns into \[\large \frac{8}{3} \div \frac{13}{16}\]

jimthompson5910 (jim_thompson5910):

now you flip the second fraction and multiply \[\large \frac{8}{3} \div \frac{13}{16}\] \[\large \frac{8}{3} \times \frac{16}{13}\]

OpenStudy (anonymous):

128/39

jimthompson5910 (jim_thompson5910):

very good \[\large \frac{8}{3} \div \left(\frac{1}{4} + \frac{9}{16}\right)\] \[\large \frac{8}{3} \div \frac{13}{16}\] \[\large \frac{8}{3} \times \frac{16}{13}\] \[\large \frac{8*16}{3*13}\] \[\large \frac{128}{39}\]

OpenStudy (anonymous):

dont we divide now

jimthompson5910 (jim_thompson5910):

you can leave it as an improper fraction or you can convert that to a mixed number either way works

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome